# Deployability-Centric Infrastructure-as-Code Generation: Fail, Learn, Refine, and Succeed through LLM-Empowered DevOps Simulation

### v2

@@ 4.3. The Iterative Feedback
Previous research have shown the effectiveness of feedback mechanisms in code generation tasks (chen2023teaching; yang2022generating; huang2022large; zhou2024bridging). IaC deployment inherently involves continuous integration and development with cycles of validation and refinement (sokolowski2022infrastructure; aws-best-practices). By integrating a feedback loop mechanism into our framework, we simulate this real-world IaC development workflow while treating LLMs as “human developers”, evaluating not only an LLM’s initial generation capability, but also its error-resolution capacity. Our feedback design implements a progressive assistance approach that mirrors established DevOps practices for IaC development. This iterative feedback reflects how cloud engineers typically engage with increasingly specific guidance when resolving deployment errors.
@@
Detailed Feedback. It offers more specific guidance by including complete error messages returned by the validation tools (yamllint, cfn-lint, and AWS SDK). For instance, instead of simply indicating a deployment error, the LLM receives the full error message: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [KeyName] must have values.” This level of feedback approximates the detailed diagnostic information that developers typically encounter when working with IaC testing tools (hasan2020testing) and tests the LLM’s ability to interpret technical error information. Considering the experimental efficiency and practicability, we constrain feedback loops to a maximum of six iterations per stage, with two attempts for general feedback and four attempts for detailed feedback. The asymmetrical allocation strategy is consistent with the empirical observations and experimental design in previous work (palavalli2024using).
@@ 5. Experimental Settings
We select six representative leading LLMs, GPT-4o, GPT-o3-mini, Claude-3.5, Claude-3.7, DeepSeek-R1, and DeepSeek-V3. Those models cover both open-sourced and closed-sourced ecosystems, broadly demonstrating competitive performance in general code generation tasks (chen2021evaluating; sobo2025evaluating; islam2025codesim; guo2024deepseek) and have significant industry adoption (pan2024textit; deepseek-share). All evaluations are conducted through the APIs from official providers ensuring reproducibility and standardization. All models are evaluated with consistent temperature settings of 0 to maximize deterministic outputs and configured with the 8,000 maximum output token limit to accommodate complex IaC templates. We run those models on the benchmark introduced in Section 3.
@@
Evaluation Metrics. The unbiased version of pass@k (chen2021evaluating) is a standard metric used by current benchmarks (chen2021evaluating; austin2021program; liu2023your), which indicates the probability that at least one of the $k$ generated samples yields a correct solution. However, it does not effectively capture the performance of iterative feedback. To address this limitation, we introduce passItr@n, where $n$ represents the maximum number of iterations allowed. PassItr@n indicates the success rate that the IaC template is correctly generated (able to be deployed successfully) before or at iteration $n$.
@@ 6.1. RQ1: How effective is IaCGen in IaC template generation?
Table 2 presents the performance of our framework on the DPIaC-Eval benchmark as the iteration increases. The initial low success rates (passItr@1) align with the previous finding (kon2024iac), with an average score of approximately 24.7% on the first attempt of all models. This demonstrates the inherent challenge of LLM-based IaC template generation compared to general-purpose code generation tasks, where the average score of the models is 95.6% in HumanEval benchmark (chen2021evaluating).
@@
Effectiveness of Conversation History. Unlike previous research (kon2024iac; palavalli2024using; ragothaman2024optimizing; madaan2023self) which isolates each feedback iteration, our approach preserves the entire conversation history for each iteration. To evaluate the effectiveness of our iterative feedback mechanism that provides LLMs with the entire conversation history, we conduct an ablation study that compares it to the baseline approach, which only feeds the system prompt, original prompt, error message, and incorrect IaC template to the LLM. We select Claude-3.5, the best performing model, for this study. As illustrated in Fig. 8, both approaches demonstrate comparable performance during the first two iterations. However, our approach establishes a consistent advantage starting with the third iteration. Our method achieves a deployable IaC template in an average of 4.55 iterations, while the baseline approach required an average of 5.41 iterations, which is an 15.9% reduction in required iterations.
@@
Generalizability of IaCGen. IaCGen’s modular design and the cross-language capabilities of LLMs enables seamless adaptation to different IaC technologies. Practitioners can adapt IaCGen to alternative IaC languages and cloud service providers by substituting the corresponding linters (for format and syntax validation) and deployment components (for deployment verification). To evaluate IaCGen’s generalizability quantitatively, we conduct experiments on the IaC-Eval benchmark (kon2024iac), which contains Terraform templates, with updates to the linter. Due to fundamental architectural differences between CloudFormation and Terraform (discussed in Section 2.2), we focus on syntax validation following the established methodology of IaC-Eval, using Claude-3.5 based on its superior performance in our primary experiments. IaCGen demonstrates strong generalizability to Terraform, achieving 79.7% passItr@1 and 100% passItr@7 accuracy with an average of 1.58 iterations. For comparison, IaCGen requires 2.53 iterations on average to achieve 100% passItr@5 accuracy on DPIaC-Eval’s CloudFormation templates. This difference reflects the higher complexity of DPIaC-Eval templates, evidenced by their higher average lines of code and resource complexity.
@@ 6.3. RQ3: How effective is human-in-the-loop feedback strategy for IaCGen?
Human-in-the-loop feedback has emerged as a crucial strategy across machine learning and LLM domains (mosqueira2023human; zanzotto2019human; xiao2023llm), demonstrating significant improvements in complex scenarios especially in SE tasks. Enlightened by this, we incorporate human feedback into our framework to explore its capability on tackling unique challenges of IaC template generation. This level of feedback simulates the guidance that a senior cloud engineer might provide to a junior developer during code review, testing whether LLMs can effectively utilize and incorporate direct expert instructions.
@@ 6.4. RQ4: How trustworthy are the IaC templates generated by IaCGen?
User Intent Matching. User requirements for IaC templates are often ambiguously specified due to the inherent complexity and technical nature of cloud infrastructure. To evaluate how well the generated IaC templates align with the intended infrastructure requirements, we employ an intent match evaluation. Specifically, based on the DPIaC-Eval benchmark, a cloud engineering expert formulates intent specifications using Checkov’s custom policy (checkov-custom-policy), comprising two elements:
@@
Security. Security of IaC templates represents a critical concern in cloud computing, as insecure configurations can lead to data breaches, unauthorized access, and compliance violations (drosos2024your). To evaluate the security posture of LLM-generated IaC templates that passed deployability evaluation, we perform the security analysis by employing Checkov (checkov), an popular open source policy-as-code tool for IaC compliance scans (minna2025analyzing) with common industry standards such as the Center for Internet Security (cis) and AWS Foundations Benchmark (cis-aws). We define three distinct compliance metrics:
@@ 7.1. Security Policy Violations in IaC Templates
As demonstrated in our RQ4 evaluation, current models exhibit limited sensitivity to security concerns, potentially leading to vulnerabilities such as data breaches and service outages. Our security analysis identified 70 distinct policy violations across all generated IaC templates, with a severity distribution of 36 low-severity, 17 medium-severity, 7 high-severity, and 10 informational-level violations according to Prisma Cloud classifications (prisma_cloud). We further analyze the top 10 most frequently violated policies, revealing one medium-severity, seven low-severity, and two informational-level violations. This distribution indicates that LLMs not only fail to implement security best practices but also potentially introduce critical vulnerabilities.
The medium-severity policy: “AWS SNS topic has SSE disabled”, reflects one systematic security oversight where LLMs fail to enable server-side encryption for message queuing services (medium_severity_policy) in generated templates. This omission exposes sensitive data in transit and at rest, violates data protection principles, and potentially compromises confidential information processed through messaging systems. In addition, violation of the low-severity policy: “AWS security groups allow ingress from 0.0.0.0/0 to port 80”, can expose AWS resources to potential security threats (low_severity_policy). Notably, low-severity does not mean it necessarily causes low security impact (dahabiyeh2021ignorance; low_severity_impact). Although classified as low-severity, this configuration creates a direct attack vector by allowing traffic from all IP addresses to the web service port. Such misconfiguration can facilitate reconnaissance attacks, denial-of-service attempts, and unauthorized access to application endpoints (mathew2014study).
The security policy violations could stem from the problematic training data. Analysis of publicly available IaC templates reveals that many contain insecure default configurations, such as security groups configured with “0.0.0.0/0” IP ranges as placeholder values (public_problematic_template_security_group), or SNS topics deployed without encryption keys for operational simplicity (public_problematic_template_SNS). LLMs appear to be pre-trained and perpetuate these problematic patterns without understanding their security implications.
@@ 7.2. Human Feedback: Automation Potential
Even though the Claude models can achieve satisfying results (94.15% passItr@15) without human intervention, results show human feedback can enhance the overall performance. To explore the possibility on automation human feedback, we conduct a post-hoc analysis. We surprisingly find that most of human feedback could be derived from publicly available sources, such as AWS documentation (aws-doc), Stack Overflow (stackoverflow) discussions, and technical blogs. Three representative cases demonstrate how human feedback can be obtained from crowd knowledge:
S3 Bucket Naming Conflicts. GPT models commonly assign generic bucket names, causing deployment failures due to S3’s global uniqueness requirement. The original human feedback is: “S3 bucket names must be globally unique across all AWS accounts. Consider adding a unique identifier such as: random string, organization prefix, timestamp, or environment identifier to ensure uniqueness”, which guides the model to generate unique names following established naming conventions. This principle directly corresponds to the best practices documented in technical resources (bucket_naming_solution; bucket_naming_guide), which explicitly stipulate name uniqueness (“Bucket names must be unique across all AWS accounts in all of the AWS”) and solutions (“Include a Randomized Suffix or Prefix”).
Outdated AMI References. LLMs frequently provide invalid AMI identifiers for EC2 instances probably caused by biased and outdated training data, as AMI changes and is different between regions. Human feedback directs models to refer to AWS Systems Manager Parameter Store for dynamic AMI resolution. This approach follows established instructions documented in official AWS blogs (ami_solution), which demonstrate programmatic AMI lookup using the latest Amazon Linux AMI parameters (“Querying for the latest AMI using public parameters”).
Service-Specific Configuration Constraints. Complex AWS service interdependencies often result in attribute-level errors, such as the constraint “DbSubnetGroupName should not be specified for read replicas created in the same region as the master.” Human feedback resolves this by explicitly instructing LLMs to remove the conflicting attribute. Similarly, the solution is directly related to the AWS documentation (dbinstance_solution), which explicitly describes this constraint in the DbSubnetGroupName specification (“The specified DB subnet group must be in the same AWS Region…”).
@@ 8.1. Feedback Mechanism of LLMs
Feedback mechanism has been widely adopted to enhance LLM outputs in SE tasks. For instance, Chen et al. (chen2023teaching) used natural language feedback to refine code completions, Huang et al. (huang2022large) introduced feedback signals during model training, and Yang et al. (yang2022generating) leveraged feedback for iterative output correction. Despite these advances, the use of feedback in IaC generation remains underexplored, particularly when considering the need for executable and deployable outputs.
Several studies have incorporated iterative feedback into LLM-based code generation (zhang2023repocoder; bi2024iterative; bouzenia2025you). However, these approaches typically provide only immediate error messages and problematic code snippets. Due to IaC’s declarative nature and complex resource interdependencies, addressing one error often introduces new issues elsewhere, as fixes frequently require coordinated changes across multiple template sections. Thus, in this study, we provide the complete conversation history to the LLM, including all previous errors and attempted solutions, enabling in-context learning (dong2022survey) and a more informed decision-making that avoids repeating previous mistakes.
@@ 8.2. LLMs for IaC Generation
Despite the growing importance of IaC in modern cloud development, research on LLMs’ capability to generate reliable infrastructure code remains limited. Kon et al. (kon2024iac) introduced the first benchmark for evaluating LLM-generated IaC, focusing primarily on syntactic validity and intent compliance across Terraform templates. They found LLMs achieved less than 20% success rates, which is substantially lower than general code generation tasks. Ragothaman (ragothaman2024optimizing) attempted to address deployability concerns but provided limited technical details about their validation process and relied heavily on manual testing procedures. While they identified the potential value of incorporating error feedback from failed templates into prompts or retrieval-augmented generation systems, they did not implement or evaluate such mechanisms. Similarly, Palavalli et al. (palavalli2024using) examined feedback loops for IaC generation, discovering that feedback effectiveness diminishes significantly after five iterations. However, their evaluation remained limited to syntax validation, and their feedback strategy only provided incorrect IaC template and basic error information.

### v3

@@ 4.3. The Iterative Feedback
Previous research have shown the effectiveness of feedback mechanisms in code generation tasks (Chen et al., 2023; Yang et al., 2022; Huang et al., 2022; Zhou et al., 2024). IaC deployment inherently involves continuous integration and development with cycles of validation and refinement (Sokolowski, 2022; AWS, 2025b). By integrating a feedback loop mechanism into our framework, we simulate this real-world IaC development workflow while treating LLMs as “human developers”, evaluating not only an LLM’s initial generation capability, but also its error-resolution capacity. Our feedback design implements a progressive assistance approach that mirrors established DevOps practices for IaC development. This iterative feedback reflects how cloud engineers typically engage with increasingly specific guidance when resolving deployment errors.
@@
Detailed Feedback. It offers more specific guidance by including complete error messages returned by the validation tools (yamllint, cfn-lint, and AWS SDK). For instance, instead of simply indicating a deployment error, the LLM receives the full error message: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [KeyName] must have values.” This level of feedback approximates the detailed diagnostic information that developers typically encounter when working with IaC testing tools (Hasan et al., 2020) and tests the LLM’s ability to interpret technical error information. Considering the experimental efficiency and practicability, we constrain feedback loops to a maximum of six iterations per stage, with two attempts for general feedback and four attempts for detailed feedback. The asymmetrical allocation strategy is consistent with the empirical observations and experimental design in previous work (Palavalli and Santolucito, 2024).
@@ 5. Experimental Settings
We select six representative leading LLMs, GPT-4o, GPT-o3-mini, Claude-3.5, Claude-3.7, DeepSeek-R1, and DeepSeek-V3. Those models cover both open-sourced and closed-sourced ecosystems, broadly demonstrating competitive performance in general code generation tasks (Chen et al., 2021; Sobo et al., 2025; Islam et al., 2025; Guo et al., 2024; Tao et al., 2025; Gong et al., 2025) and have significant industry adoption (Pan et al., 2024b; Kumar, 2025). All evaluations are conducted through the APIs from official providers ensuring reproducibility and standardization. All models are evaluated with consistent temperature settings of 0 to maximize deterministic outputs and configured with the 8,000 maximum output token limit to accommodate complex IaC templates. We run those models on the benchmark introduced in Section 3.
@@
Evaluation Metrics. The unbiased version of pass@k (Chen et al., 2021) is a standard metric used by current benchmarks (Chen et al., 2021; Austin et al., 2021; Liu et al., 2023), which indicates the probability that at least one of the $k$ generated samples yields a correct solution. However, it does not effectively capture the performance of iterative feedback. To address this limitation, we introduce passItr@n, where $n$ represents the maximum number of iterations allowed. PassItr@n indicates the success rate that the IaC template is correctly generated (able to be deployed successfully) before or at iteration $n$.
@@ 6.1. RQ1: How effective is IaCGen in IaC template generation?
Table 2 presents the performance of our framework on the DPIaC-Eval benchmark as the iteration increases. The initial low success rates (passItr@1) align with the previous finding (Kon et al., 2024), with an average score of approximately 24.7% on the first attempt of all models. This demonstrates the inherent challenge of LLM-based IaC template generation compared to general-purpose code generation tasks, where the average score of the models is 95.6% in HumanEval benchmark (Chen et al., 2021).
@@
Effectiveness of Conversation History. Unlike previous research (Kon et al., 2024; Palavalli and Santolucito, 2024; Ragothaman and Udayakumar, 2024; Madaan et al., 2023) which isolates each feedback iteration, our approach preserves the entire conversation history for each iteration. To evaluate the effectiveness of our iterative feedback mechanism that provides LLMs with the entire conversation history, we conduct an ablation study that compares it to the baseline approach, which only feeds the system prompt, original prompt, error message, and incorrect IaC template to the LLM. We select Claude-3.5, the best performing model, for this study. As illustrated in Fig. 8, both approaches demonstrate comparable performance during the first two iterations. However, our approach establishes a consistent advantage starting with the third iteration. Our method achieves a deployable IaC template in an average of 4.55 iterations, while the baseline approach required an average of 5.41 iterations, which is an 15.9% reduction in required iterations.
@@
Generalizability of IaCGen. IaCGen’s modular design and the cross-language capabilities of LLMs enables seamless adaptation to different IaC technologies. Practitioners can adapt IaCGen to alternative IaC languages and cloud service providers by substituting the corresponding linters (for format and syntax validation) and deployment components (for deployment verification). To evaluate IaCGen’s generalizability quantitatively, we conduct experiments on the IaC-Eval benchmark (Kon et al., 2024), which contains Terraform templates, with updates to the linter. Due to fundamental architectural differences between CloudFormation and Terraform (discussed in Section 2.2), we focus on syntax validation following the established methodology of IaC-Eval, using Claude-3.5 based on its superior performance in our primary experiments. IaCGen demonstrates strong generalizability to Terraform, achieving 79.7% passItr@1 and 100% passItr@7 accuracy with an average of 1.58 iterations. For comparison, IaCGen requires 2.53 iterations on average to achieve 100% passItr@5 accuracy on DPIaC-Eval’s CloudFormation templates. This difference reflects the higher complexity of DPIaC-Eval templates, evidenced by their higher average lines of code and resource complexity.
@@ 6.3. RQ3: How effective is human-in-the-loop feedback strategy for IaCGen?
Human-in-the-loop feedback has emerged as a crucial strategy across machine learning and LLM domains (Mosqueira-Rey et al., 2023; Zanzotto, 2019; Xiao and Wang, 2023), demonstrating significant improvements in complex scenarios especially in SE tasks. Enlightened by this, we incorporate human feedback into our framework to explore its capability on tackling unique challenges of IaC template generation. This level of feedback simulates the guidance that a senior cloud engineer might provide to a junior developer during code review, testing whether LLMs can effectively utilize and incorporate direct expert instructions.
@@ 6.4. RQ4: How trustworthy are the IaC templates generated by IaCGen?
User Intent Matching. User requirements for IaC templates are often ambiguously specified due to the inherent complexity and technical nature of cloud infrastructure. To evaluate how well the generated IaC templates align with the intended infrastructure requirements, we employ an intent match evaluation. Specifically, based on the DPIaC-Eval benchmark, a cloud engineering expert formulates intent specifications using Checkov’s custom policy (Checkov, n.d.), comprising two elements:
@@
Security. Security of IaC templates represents a critical concern in cloud computing, as insecure configurations can lead to data breaches, unauthorized access, and compliance violations (Drosos et al., 2024). To evaluate the security posture of LLM-generated IaC templates that passed deployability evaluation, we perform the security analysis by employing Checkov (Nimrod et al., 2025), an popular open source policy-as-code tool for IaC compliance scans (Minna et al., 2025) with common industry standards such as the Center for Internet Security (for Internet Security, n.d.) and AWS Foundations Benchmark (for Internet Security, 2025). We define three distinct compliance metrics:
@@ 7.1. Security Policy Violations in IaC Templates
As demonstrated in our RQ4 evaluation, current models exhibit limited sensitivity to security concerns, potentially leading to vulnerabilities such as data breaches and service outages. Our security analysis identified 70 distinct policy violations across all generated IaC templates, with a severity distribution of 36 low-severity, 17 medium-severity, 7 high-severity, and 10 informational-level violations according to Prisma Cloud classifications (Cloud, n.d.). We further analyze the top 10 most frequently violated policies, revealing one medium-severity, seven low-severity, and two informational-level violations. This distribution indicates that LLMs not only fail to implement security best practices but also potentially introduce critical vulnerabilities.
The medium-severity policy: “AWS SNS topic has SSE disabled”, reflects one systematic security oversight where LLMs fail to enable server-side encryption for message queuing services (Cloud, 2025b) in generated templates. This omission exposes sensitive data in transit and at rest, violates data protection principles, and potentially compromises confidential information processed through messaging systems. In addition, violation of the low-severity policy: “AWS security groups allow ingress from 0.0.0.0/0 to port 80”, can expose AWS resources to potential security threats (Cloud, 2025a). Notably, low-severity does not mean it necessarily causes low security impact (Dahabiyeh et al., 2021; Carpenter, 2024). Although classified as low-severity, this configuration creates a direct attack vector by allowing traffic from all IP addresses to the web service port. Such misconfiguration can facilitate reconnaissance attacks, denial-of-service attempts, and unauthorized access to application endpoints (Mathew et al., 2014).
The security policy violations could stem from the problematic training data. Analysis of publicly available IaC templates reveals that many contain insecure default configurations, such as security groups configured with “0.0.0.0/0” IP ranges as placeholder values (Casas, 2025), or SNS topics deployed without encryption keys for operational simplicity (Beard, 2024). LLMs appear to be pre-trained and perpetuate these problematic patterns without understanding their security implications.
@@ 7.2. Human Feedback: Automation Potential
Even though the Claude models can achieve satisfying results (94.15% passItr@15) without human intervention, results show human feedback can enhance the overall performance. To explore the possibility on automation human feedback, we conduct a post-hoc analysis. We surprisingly find that most of human feedback could be derived from publicly available sources, such as AWS documentation (AWS, 2025a), Stack Overflow (Overflow, n.d.) discussions, and technical blogs. Three representative cases demonstrate how human feedback can be obtained from crowd knowledge:
S3 Bucket Naming Conflicts. GPT models commonly assign generic bucket names, causing deployment failures due to S3’s global uniqueness requirement. The original human feedback is: “S3 bucket names must be globally unique across all AWS accounts. Consider adding a unique identifier such as: random string, organization prefix, timestamp, or environment identifier to ensure uniqueness”, which guides the model to generate unique names following established naming conventions. This principle directly corresponds to the best practices documented in technical resources (Kuma, 2024; AWS, n.d.b), which explicitly stipulate name uniqueness (“Bucket names must be unique across all AWS accounts in all of the AWS”) and solutions (“Include a Randomized Suffix or Prefix”).
Outdated AMI References. LLMs frequently provide invalid AMI identifiers for EC2 instances probably caused by biased and outdated training data, as AMI changes and is different between regions. Human feedback directs models to refer to AWS Systems Manager Parameter Store for dynamic AMI resolution. This approach follows established instructions documented in official AWS blogs (Yip, 2018), which demonstrate programmatic AMI lookup using the latest Amazon Linux AMI parameters (“Querying for the latest AMI using public parameters”).
Service-Specific Configuration Constraints. Complex AWS service interdependencies often result in attribute-level errors, such as the constraint “DbSubnetGroupName should not be specified for read replicas created in the same region as the master.” Human feedback resolves this by explicitly instructing LLMs to remove the conflicting attribute. Similarly, the solution is directly related to the AWS documentation (AWS, n.d.a), which explicitly describes this constraint in the DbSubnetGroupName specification (“The specified DB subnet group must be in the same AWS Region…”).
@@ 8.1. Feedback Mechanism of LLMs
Feedback mechanism has been widely adopted to enhance LLM outputs in SE tasks. For instance, Chen et al. (Chen et al., 2023) used natural language feedback to refine code completions, Huang et al. (Huang et al., 2022) introduced feedback signals during model training, and Yang et al. (Yang et al., 2022) leveraged feedback for iterative output correction. Despite these advances, the use of feedback in IaC generation remains underexplored, particularly when considering the need for executable and deployable outputs.
Several studies have incorporated iterative feedback into LLM-based code generation (Zhang et al., 2023; Bi et al., 2024; Bouzenia and Pradel, 2025). However, these approaches typically provide only immediate error messages and problematic code snippets. Due to IaC’s declarative nature and complex resource interdependencies, addressing one error often introduces new issues elsewhere, as fixes frequently require coordinated changes across multiple template sections. Thus, in this study, we provide the complete conversation history to the LLM, including all previous errors and attempted solutions, enabling in-context learning (Dong et al., 2022) and a more informed decision-making that avoids repeating previous mistakes.
@@ 8.2. LLMs for IaC Generation
Despite the growing importance of IaC in modern cloud development, research on LLMs’ capability to generate reliable infrastructure code remains limited. Kon et al. (Kon et al., 2024) introduced the first benchmark for evaluating LLM-generated IaC, focusing primarily on syntactic validity and intent compliance across Terraform templates. They found LLMs achieved less than 20% success rates, which is substantially lower than general code generation tasks. Ragothaman (Ragothaman and Udayakumar, 2024) attempted to address deployability concerns but provided limited technical details about their validation process and relied heavily on manual testing procedures. While they identified the potential value of incorporating error feedback from failed templates into prompts or retrieval-augmented generation systems, they did not implement or evaluate such mechanisms. Similarly, Palavalli et al. (Palavalli and Santolucito, 2024) examined feedback loops for IaC generation, discovering that feedback effectiveness diminishes significantly after five iterations. However, their evaluation remained limited to syntax validation, and their feedback strategy only provided incorrect IaC template and basic error information.